# Incorporating Procedural Fairness in Flag Submissions on Social Media Platforms

### v1

@@ 6.1.2.   Establishing Mechanisms to Track Flag Status
We echo participants’ suggestion that platforms implement a streamlined tracking system to monitor the review progress of submitted flags. Previous HCI research efforts to design moderation procedures to appeal for the reversal of unwarranted post-removals (Vaccaro et al., 2021, 2020) could serve as a blueprint for taking a user-centered approach to building such systems. Some participants expressed a desire to provide feedback on their flagging experiences (Sec.  5.2.2). This suggests that offering a chance to be heard on the entire flagging process (i.e., beyond their objections to individual flagged posts) may influence users’ fairness perceptions. Therefore, platforms could deploy a feedback system to collect recommendations for improvements and regularly evaluate the feedback received to update their current practices.
@@ 6.1.3.   Widening Flags’ Vocabulary of Complaint
Prior theoretical analysis of flagging mechanisms emphasized that flags offer “a narrow vocabulary of complaint” and do not account for the many complex reasons that people might choose to flag or let users specify their degree of concern with the flagged posts (Crawford and Gillespie, 2016). One might expect that offering a selection of more granular rule violation categories would empower users to better express their objections. In line with this, we found that a simple classification scheme enhances the perceived sense of voice more than having no classification scheme (Sec. 5.1.1). Further, offering this scheme is especially important for voice perceptions when a text box is not provided and using this scheme does not impact usability (Sec. 5.1.5). However, we found that voice perceptions do not significantly differ between simple and detailed schemes. Our findings also reveal that raising the granularity of rule violations does not significantly increase users’ perceived consistency and transparency (Sec. 5.1.1). Overall, these results suggest that although deploying simple classification schemes helps flaggers voice their objections to platform operators as compared to having no such schemes, asking users to categorize their objections this way remains fundamentally a constricted way to submit reports from users’ perspectives—and adding submenus to these schemes does not aid in addressing users’ expression and fairness needs (Sec. 5.1.1).
@@
In sum, enacting flagging design solutions that let users clarify the post’s context and offer greater flexibility than having to shoehorn complex feelings into a single category selection (Crawford and Gillespie, 2016) would enhance flaggers’ fairness perceptions.
@@ 6.2.2.   Offering Information About Flag Reviewers
Our quantitative analysis shows that fairness perceptions of flagging procedures are not influenced by information about the moderator type (a human, a bot, or no information) (Sec.  5.1.4). This result aligns with prior research findings  (Ozanne et al., 2022; Gonçalves et al., 2021; Molina and Sundar, 2022). However, our qualitative findings suggest that offering other types of information about moderators who review flags may enhance users’ perceptions of fairness. Participants suggested that disclosing moderators’ personal characteristics, e.g., information about their experiences and skills, such as the training moderators receive or their expertise in specific subjects, can foster greater trust in the flag review procedures (Sec.  5.2.1). Thus, we suggest that flagging mechanisms offer some information about flag reviewers while still respecting their privacy.
Our findings highlight the utility of involving multiple reviewers in content moderation, reflecting users’ preferences for incorporating diverse perspectives (Sec.  5.2.1). Participants suggested that using more than one moderator or combining human and bot reviewers could enhance fairness in the flag review process. This aligns with Fan and Zhang’s finding that group-based moderation improves perceptions of fairness (Fan and Zhang, 2020) and Katsaros et al.’s observation that users prefer systems that combine humans with algorithmic decisions (Katsaros et al., 2024), supporting the need for hybrid review models. Our qualitative findings, as well as prior research (Duguay et al., 2020; Pan et al., 2022), suggest that users’ preference for involving multiple moderators stems from their concerns about potential biases in decisions made by a single moderator, particularly human moderators who might be influenced by their personal values when making moderation decisions. Therefore, flagging mechanisms can enhance their fairness by instituting procedures that incorporate the perspectives of multiple flag reviewers and enacting substantive measures to prevent biases in moderators’ decision-making.
@@ 6.2.3.   Reforming Post-flag Submission Steps
Participants emphasized the need for greater transparency in the flag review process and the disclosure of information about flag submission steps. This involves revealing the review criteria, the specifics of each step in the flag review, the expected timeline, and the information visible to reviewers about flaggers (Sec.  5.2.4).
@@
Participants also considered it important that reviewers rigorously assess whether the flagged post indeed violates the inappropriate content criteria specified in the flag (Sec.  5.2.5). Thus, we recommend that platforms develop a publicly accessible protocol for handling situations in which the flagged post does not actually fall into the category (e.g., hate speech, misinformation) that flaggers choose. For instance, if the flag review determines that the flagged post is inappropriate but for a different reason than that included in the flag, flag reviewers may still sanction that post and additionally use this opportunity to inform flaggers about how to make accurate selections from the rule violation classification scheme; such educational measures may reduce the flag review burden for subsequent flags. Alternatively, flag reviewers may allow flaggers to re-classify the flag since this could increase the certainty of the review’s legitimacy.
6.3. Supporting Users with Diverse Technological Competencies and Perspectives.
@@ 6.4.2.   Encouraging Norm Compliance Among Flagged Users
Therefore, platforms should consider investing in educational measures, such as helping the authors of flagged posts better understand the community rules and how to adhere to them. As prior research shows, moderator messages that explain to the sanctioned users why their posts are removed helps improve their attitudes and future behaviors (Jhaver et al., 2019a, c). Highlighting, rewarding, and incentivizing desirable behaviors also reinforce constructive contributions (Lambert et al., 2024; Choi et al., 2024). Further, platforms may strengthen ex ante moderation measures (Grimmelmann, 2015), such as surfacing posting guidelines while a user is writing a post or using AI-based tools to warn users when their post draft is likely to be sanctioned (Jerasa and Burriss, 2024; Katsaros et al., 2022b). Such measures may reduce the burden of flagging on regular users and flag review on moderators.
Additionally, platforms might consider adopting soft moderation measures (instead of content removal) in some cases, such as putting information labels to contextualize the information in flagged posts (Morrow et al., 2022) or enacting a strike system for flagged accounts. Such measures would contribute to educating flagged users and encouraging constructive posts in the future (Singhal et al., 2023).
@@ 6.5. Limitations and Future Work
Our experiments evaluated participants’ flagging of a single inappropriate post. However, users’ behaviors may vary in their daily social media use as they encounter multiple instances of norm violations. Therefore, longitudinal analyses that examine how users interact with flagging tools in their day-to-day settings would be valuable to further inform the tradeoffs between incorporating fairness and reducing cognitive burden.
Flagging is one of a suite of moderation mechanisms available to end-users, e.g., instead of flagging, users could choose to downvote posts (Seredko and Hillman, 2024), mute or block norm-violating accounts (Geiger, 2016; Jhaver et al., 2018), engage in counter speech (Garland et al., 2020), or configure sensitivity controls for their newsfeeds (Jhaver et al., 2022, 2023). Each such mechanism offers a different set of affordances and limitations, e.g., blocking tools guarantee an instant obscuring of inappropriate content and are more free speech preserving (Jhaver and Zhang, 2023), but they have little effect on posters of that content. It would be valuable to identify the various psychological factors and social influences shaping users’ selections of moderation mechanisms to address different kinds of online harm.
@@ 7. Conclusion
This paper examines how including flag components that provide different types of information about the flag review process shape users’ attitudes toward flagging. Our analysis shows that users desire flag mechanisms to support greater expressivity, timely notifications of flag review updates, increased visibility into flag review procedures and reviewers, and stronger protections against flag abuse. We discuss how innovations in flagging systems, such as improving their accessibility and offering support for highlighting the severity of rule violations, could better support end-users’ fairness demands. This investigation demonstrates how the design and policy choices made in the implementation of flagging infrastructures deeply shape users’ daily experiences of social media use and address (or fail to address) their vital needs to combat content-based harms. We call for future studies to deploy similar user-centered approaches and social justice orientations to improve the current practices of platform governance.
@@ Appendix A Survey Sample Description
Appendix B Interaction effects
Table 13 below presents the GLM analysis results, which indicate how different combinations of flagging components interact, affecting perceived consistency, transparency, and voice. Next, we present Table 14, which details our results on interaction effects of the classification scheme and a text box on perceived voice. We comment on these results in sec. 5.1.5.

### v2

@@ 6.1.2.   Establishing Mechanisms to Track Flag Status
We echo participants’ suggestion that platforms implement a streamlined tracking system to monitor the review progress of submitted flags. Previous HCI research efforts to design for contestability in moderation procedures (Vaccaro et al., 2021, 2020) could serve as a blueprint for taking a user-centered approach to building such systems. Some participants expressed a desire to provide feedback on their flagging experiences (Sec.  5.2.2). This suggests that offering a chance to be heard on the entire flagging process (i.e., beyond their objections to individual flagged posts) may influence users’ fairness perceptions. Therefore, platforms could deploy a feedback system to collect recommendations for improvements and regularly evaluate the feedback received to update their current practices.
@@ 6.1.3.   Widening Flags’ Vocabulary of Complaint
Prior theoretical analysis of flagging mechanisms emphasized that flags offer “a narrow vocabulary of complaint” and do not account for the many complex reasons that people might choose to flag or let users specify their degree of concern with the flagged posts (Crawford and Gillespie, 2016). Our quantitative analysis did not find a relationship between the granularity of classification schemes and fairness perceptions (Sec. 5.1.1). However, we found that offering a classification scheme (either simple or detailed) enhances voice perceptions when a text box is not provided (Table 11) and providing this scheme does not impact usability (Sec. 5.1.5). Overall, these results suggest that asking users to categorize their objections using a classification scheme remains fundamentally a constricted way to submit reports from users’ perspectives—and adding submenus to these schemes does not aid in addressing users’ expression and fairness needs (Sec. 5.1.1).
@@
While users desire expressive flagging options, it is crucial that platforms provide such options only when they can effectively incorporate that expression. For instance, allowing flaggers to provide free-text responses may enhance their sense of having a voice (perceived fairness) during flag submission, but it does not improve actual procedural fairness if the flag reviewers lack the ability to process that input when making its moderation decisions.
In sum, enacting flagging solutions that let users clarify the post’s context and offer greater flexibility than having to shoehorn complex feelings into a single category selection (Crawford and Gillespie, 2016) would enhance flaggers’ fairness perceptions.
@@ 6.2.2.   Offering Information About Flag Reviewers
Our quantitative analysis shows that fairness perceptions of flagging procedures are not influenced by information about the moderator type (a human, a bot, or no information) (Sec.  5.1.4). This result aligns with prior research findings  (Ozanne et al., 2022; Gonçalves et al., 2021; Molina and Sundar, 2022). However, our qualitative findings suggest that offering other types of information about moderators who review flags may enhance users’ perceptions of fairness. Participants suggested that disclosing moderators’ personal characteristics, e.g., information about their experiences and skills, such as the training moderators receive or their expertise in specific subjects, can foster greater trust in the flag review procedures (Sec.  5.2.1). Thus, we suggest that flagging mechanisms offer some information about flag reviewers while still respecting their privacy. Such information could also enhance moderation fairness from the perspective of flagged users.
Our findings highlight the utility of involving multiple reviewers in content moderation, reflecting users’ preferences for incorporating diverse perspectives (Sec.  5.2.1). Participants suggested that using more than one moderator or combining human and bot reviewers could enhance fairness in the flag review process. This aligns with Fan and Zhang’s finding that group-based moderation improves perceptions of fairness (Fan and Zhang, 2020) and Katsaros et al.’s observation that users prefer systems that combine humans with algorithmic decisions (Katsaros et al., 2024), supporting the need for hybrid review models. Our qualitative findings, as well as prior research (Duguay et al., 2020; Pan et al., 2022), suggest that users’ preference for involving multiple moderators stems from their concerns about potential biases in decisions made by a single moderator, particularly human moderators who might be influenced by their personal values when making moderation decisions (Sec.  5.2.1). Therefore, flagging mechanisms can enhance their fairness by instituting procedures that incorporate the perspectives of multiple flag reviewers and enacting substantive measures to prevent biases in moderators’ decision-making.
@@ 6.2.3.   Reforming Post-flag Submission Steps
Participants emphasized the need for greater transparency in the flag review process and the disclosure of information about post-flag submission steps. This involves revealing the review criteria, the specifics of each step in the flag review, the expected timeline, and the information visible to reviewers about flaggers (Sec.  5.2.4).
@@
6.3. Supporting Flaggers with Different Technological Competencies and Diverse Perspectives.
@@ 6.4.2.   Encouraging Norm Compliance Among Flagged Users
Therefore, platforms should consider investing in educational measures, such as helping the authors of flagged posts better understand the community rules and how to adhere to them. As prior research shows, moderator messages that explain to the sanctioned users why their posts are removed helps improve their attitudes and future behaviors (Jhaver et al., 2019a, c). Highlighting, rewarding, and incentivizing desirable behaviors also reinforce constructive contributions (Lambert et al., 2024; Choi et al., 2024). Further, platforms may strengthen ex ante moderation measures (Grimmelmann, 2015), such as surfacing posting guidelines while a user is writing a post or using AI-based tools to warn users when their post draft is likely to be sanctioned (Jerasa and Burriss, 2024; Katsaros et al., 2022b). In line with findings from Zhang et al. (2023), some of our respondents preferred directly communicating with rule-violators over reporting them to persuade norm compliance. Such measures may reduce the burden of flagging on regular users and flag review on moderators.
@@ 6.5. Limitations and Future Work
We captured our participants’ demographic details directly from Lucid’s pre-collected data, which does not allow reporting of non-binary gender identities. Future studies using Lucid Theorem can capture this information using a separate question in the survey. While we undertook pre-processing steps to discard inappropriate participant responses, implementing attention checks in our survey questionnaire would have further improved the validity of this study.
Our experiments evaluated participants’ flagging of a single inappropriate post. However, users’ behaviors may vary in their daily social media use as they encounter multiple instances of norm violations. Additionally, while the use of highly offensive stimuli in our survey creates a strong setting for evaluating flag mechanisms, less extreme content might shape flagging needs and actions differently. Therefore, longitudinal or in situ analyses that examine how users interact with flagging tools in their day-to-day settings would be valuable to further inform the tradeoffs between incorporating fairness and reducing cognitive burden.
@@ 7. Conclusion
This paper examines how flag components that provide different types of information about the flag review process shape users’ attitudes toward flagging. Our analysis shows that including posting guidelines in flag designs enhances users’ transparency perceptions and offering a text box improves their voice perceptions. We found that users desire flag mechanisms to support greater expressivity, timely notifications of flag review updates, increased visibility into flag review procedures and reviewers, and stronger protections against flag abuse. We discuss how innovations in flagging systems, such as improving their accessibility and offering support for highlighting the severity of rule violations, could better support end-users’ fairness demands. This investigation demonstrates how the design and policy choices made in the implementation of flagging infrastructures deeply shape users’ daily experiences of social media use and address (or fail to address) their vital needs to combat content-based harms. We call for future studies to deploy similar user-centered approaches and social justice orientations to improve the current practices of platform governance.
@@ Appendix A Survey Sample Description
Appendix B Fairness Perceptions for Different Flagged Posts
Table 9 shows variations in fairness perceptions for different selections of inappropriate posts to flag by survey respondents.
Table 9. The mean values of perceived consistency, transparency, and voice for each of the rule violation examples that participants choose to flag in the survey. Standard deviation (SD) values are shown in brackets.
Category N Mean (SD)
Consistency Transparency Voice
Mexicans come from an uncivilized … 621 5.57 (1.45) 5.33 (1.31) 5.31 (1.49)
@Sean11 I hate all you F*** … 272 5.50 (1.52) 5.39 (1.29) 5.33 (1.43)
Drinking bleach has been scientifically … 694 5.68 (1.37) 5.42 (1.36) 5.35 (1.56)
Today is a great day! I ate 723 calories … 165 5.09 (1.62) 4.98 (1.64) 5.01 (1.74)
Download the software for credit card … 178 5.34 (1.57) 5.01 (1.49) 5.00 (1.60)
Omg just got tons of Bucks from here! … 134 5.24 (1.56) 5.00 (1.41) 4.86 (1.68)
I have a masturbation video of @Janny12 … 872 5.60 (1.39) 5.34 (1.37) 5.43 (1.51)
Total 2,936 5.54 (1.45) 5.30 (1.38) 5.30 (1.55)
Appendix C Additional Analyses
Table 10 below presents the GLM analysis results, which indicate how different combinations of flagging components interact, affecting perceived consistency, transparency, and voice. Next, we present Table 11, which details our results on interaction effects of the classification scheme and a text box on perceived voice. Tables 13 and 13 summarize our results on how different flagging components influence usability. We comment on these results in sec. 5.1.5.
@@ Appendix C Additional Analyses
Components SS df MS f/t p
Classification 8.14 2 4.07 1.31 .27
Guidelines 17.63 2 8.81 2.83 .06
Text box - 2933.87 - -3.03 .01
Moderator 4.01 2 2.00 0.64 .53
Components SS df MS f/t p
Classification .91 2 .46 .19 .83
Guidelines .12 2 .06 .03 .98
Text box - 2930.17 - -1.07 .14
Moderator 8.40 2 4.20 1.77 .17